# RLHS: Mitigating Misalignment in RLHF with Hindsight Simulation

### v2

@@ Abstract
While Reinforcement Learning from Human Feedback (RLHF) has shown promise in aligning generative AI, we present empirical evidence that it can also cause severe, systematic misalignment. We hypothesize that this stems from evaluator feedback depending on downstream outcome predictions (foresight) that can be influenced by the AI’s output, inducing Goodhart’s law dynamics. Conversely, our theoretical analysis shows that conditioning evaluator feedback on downstream observations (hindsight) inhibits this effect by decoupling the alignment signal from potentially compromised predictions—crucially, the result holds even if the observed outcomes are sampled from the AI’s own world model. Building on this insight, we introduce Reinforcement Learning from Hindsight Simulation (RLHS), which presents plausible simulated outcomes to evaluators before eliciting feedback. We demonstrate RLHS on online (PPO) and offline (DPO) large language model fine-tuning, obtaining superior alignment over RLHF in controlled consultancy-type experiments and user studies. We evaluate post-hoc on the TruthfulQA benchmark and find that, even after single-task fine-tuning, both RLHF misalignment and RLHS alignment carry over to substantially different settings.
Machine Learning, ICML
@@ 1 Introduction
Aligning artificial intelligence (AI) systems with human values and goals is crucial to ensuring their behavior is helpful, honest, and trustworthy. Eliciting human feedback is a widely-used alignment strategy (Leike et al., 2018), with successful applications to, e.g., training AI assistants (Glaese et al., 2022; Touvron et al., 2023; Anthropic, 2023; Achiam et al., 2023). In particular, Reinforcement Learning from Human Feedback (RLHF) (Christiano et al., 2017; Ziegler et al., 2019; Ouyang et al., 2022; Stiennon et al., 2020) leverages human feedback to fine-tune and align foundation models (FMs). While RLHF has shown promise in aligning models with human preferences, it relies predominantly on immediate assessments of isolated interactions, which may not accurately account for their downstream outcomes. Inaccurate user or evaluator feedback can misguide the model’s behavior and undermine the alignment process. For example, human evaluators could misjudge an interaction on the spot, due to limited knowledge (Casper et al., 2023; Lang et al., 2024) or cognitive resources  (Pandey et al., 2022; Chmielewski & Kucker, 2020), leading to incomplete or misinformed feedback. A recent study has theorized that partial observability of an AI assistant’s task execution during human–AI interaction can lead RLHF to learn deceptive behaviors (Lang et al., 2024).
@@
Figure 1: RLHF can incentivize AI systems to provide inaccurate or deceptive information to their users, prioritizing positive on-the-spot feedback and neglecting long-term consequences. For example, a customer may prefer to hear good news while shopping but will ultimately be disappointed (and objectively worse off) if stuck with an ill-informed purchase. The proposed RLHS scheme removes this incentive by simulating the interaction’s downstream outcome before eliciting feedback. The resulting fine-tuned AI models show superior alignment with users’ underlying utility.
@@
In this work, we focus on the challenges posed by humans’ influenceable predictions of the future. In many settings, the utility provided by an AI system to a human user (and similarly its “helpfulness” and “harmlessness”, which RLHF evaluators are typically asked to assess) is not an intrinsic property of the outputs that it generates but rather a function of their real-world consequences, brought about by the user’s real-world decisions upon consuming said outputs. We contend that rewarding an AI system to improve users or evaluators’ in-the-moment assessments of interactions creates a pernicious Goodhart’s law dynamic: such an incentive will favor outputs that induce unrealistically optimistic expectations in users—while at best these may be innocuous, at worst they can lead users to make poor choices resulting in degraded outcomes.
@@
We provide substantial empirical evidence that indeed this phenomenon can arise even in simple settings: we find that immediate human feedback elicited at the end of the human–AI interaction frequently misrepresents true utility in consultancy-type interactions, and, when used as a proxy for it in RLHF fine-tuning, it systematically drives misalignment with human goals (Fig. 1, top). Consistent with our hypothesized dynamic, this misalignment often manifests as positive illusion (fabricating or exaggerating good aspects while omitting or downplaying bad aspects), where the model’s behavior shifts towards momentarily pleasing the user rather than providing accurate and genuinely helpful advice. This consistently leads users to make ill-informed decisions whose poor downstream outcomes contrast starkly with their high satisfaction rating at the end of the interaction.
@@
To address these open challenges, we propose to leverage hindsight as a simple but effective misalignment mitigation mechanism, in which evaluators experience the downstream outcomes of an interaction before being asked for feedback on the model. We provide both theoretical analysis and extensive empirical studies to show the efficacy of hindsight in significantly reducing misalignment of RLHF-trained models. To circumvent the material and ethical difficulties in exposing real people to real consequences, we introduce a novel alignment algorithm called Reinforcement Learning from Hindsight Simulation (RLHS), an alternative to RLHF that uses an AI world model during training to simulate likely human decisions and their downstream outcomes. Despite lacking any privileged information on real outcomes, the method effectively relieves evaluators of the need to predict the future consequences of an AI output—thus removing the AI incentive to influence these predictions.
Our key finding is that equipping evaluators with the benefit of (simulated) hindsight—and thereby relieving them of the burden of prediction—significantly reduces model misalignment after fine-tuning. Intuitively, even if the world model generating outcomes contains inaccuracies, these are not dependent on the AI output, and therefore the AI agent has no incentive to manipulate them. We evaluate hindsight simulation with both offline and online preference optimization approaches, including direct preference optimization (DPO) (Rafailov et al., 2024) and proximal policy optimization (PPO) (Schulman et al., 2017) and find that it greatly improves alignment in both training paradigms. We validate these results through an online user study, in which RLHS consistently improved the objective utility and subjective satisfaction of our human participants, despite being trained with only simulated hindsight feedback. Our comparative findings demonstrate that RLHS significantly outperforms non-hindsight methods—specifically Reinforcement Learning from AI Feedback (RLAIF), which similarly uses AI generation as a proxy for human feedback, and has been shown to produce results similar to RLHF (Bai et al., 2022b; Lee et al., 2023). Finally, a post-hoc evaluation on the TruthfulQA benchmark indicates that, even after single-task fine-tuning, both RLHF misalignment and RLHS alignment can persist in substantially different settings.
2 Background and Preliminaries
Human Decision-Making under Uncertainty. We consider a decision problem faced by a human entity (e.g., an individual, group, or institution) under predictive uncertainty and imperfect observations. We model the a problem as a partially observable Markov decision process (POMDP) defined by a tuple $\mathcal{P}^{\textit{H}}=(\mathcal{S},\mathcal{A}^{\textit{H}},\mathcal{O}^{\textit{H}},\mathcal{T},{O}^{\textit{H}},P_{0},r,\gamma,{\theta}^{\textit{H}})$, where $\mathcal{S}$ is the set of relevant world states, $\mathcal{A}^{\textit{H}}$ is the set of available actions, $\mathcal{O}^{\textit{H}}$ is the human’s observation space, $\mathcal{T}:\mathcal{S}\times\mathcal{A}^{\textit{H}}\rightarrow\Delta(\mathcal{S})$ is the stochastic transition kernel, ${O}^{\textit{H}}:\mathcal{S}\rightarrow\Delta(\mathcal{O}^{\textit{H}})$ is the human’s observation map, $P_{0}\in\Delta(\mathcal{S})$ is the initial state distribution, $r:\mathcal{S}\times\mathcal{A}^{\textit{H}}\times\Theta^{\textit{H}}\rightarrow\mathbb{R}$ is the reward function, $\gamma\in(0,1)$ is the time discount factor, and ${\theta}^{\textit{H}}\in\Theta^{\textit{H}}$ describes the human’s intrinsic preferences. Due to partial observability of the world state ${s}\in\mathcal{S}$, the human may maintain an internal state ${z}^{\textit{H}}\in{\mathcal{Z}^{\textit{H}}}$ (e.g., a belief $b^{H}\in\Delta(\mathcal{S})$) encoding the human’s uncertain knowledge of the world state, although ${z}^{\textit{H}}$ may be thought of as a more general variable that could encode features such as the human’s emotional state or attention focus). The human may be modeled as taking actions according to a stochastic policy ${\pi}^{\textit{H}}:{\mathcal{Z}^{\textit{H}}}\rightarrow\Delta(\mathcal{A}^{\textit{H}})$.
@@ 3 Algorithm: RL from Hindsight Simulation
We begin by examining how RLHF can become misaligned due to the limitations of human foresight. We then discuss the advantages of incorporating hindsight and explain why and how it mitigates misalignment. Finally, we introduce Reinforcement Learning from Hindsight Simulation (RLHS) —a principled alignment method that employs simulated hindsight to decouple feedback from potentially AI-compromised predictions.
@@ 3.1 Hindsight Mitigates Misalignment
We consider the setting in which the human interacts with an AI up to time $k$ (interaction phase), and then proceeds to take actions (acting phase). Once the human begins to take physical actions, we are interested in the human utility, defined as a discounted sum or rewards (the reward function being unknown to the AI system). We can define the utility of any given trajectory, or equivalently, its hindsight value from the end of the horizon.
Definition 1 (Hindsight Value).
Let $\tau_{k:k+N}=({s}_{t},{o}^{\textit{H}}_{t},{a}^{\textit{H}}_{t})_{t=k}^{k+N}$ denote the trajectory of states, observations, and human actions from time $k\geq 0$ to $k+N$. The hindsight value of any such trajectory is
@@
Definition 2 (Human rollout distribution).
Let $P(\tau_{k:k+N}\mid s_{k},z^{H}_{k})$ be the probability distribution over the trajectory $\tau_{k:k+N}=(s_{t},o_{t}^{H},a_{t}^{H})_{t=k}^{k+N}$ induced by:
@@
${o}^{\textit{H}}_{t+1}$ $\sim\mathcal{O}^{\textit{H}}(\cdot\mid{s}_{t},{a}_{t}^{\textit{H}}),$ $\quad z_{t+1}^{H}$ $\sim\mathcal{T}_{z}(\cdot\mid{z}^{\textit{H}}_{t},{a}_{t}^{\textit{H}},{o}^{\textit{H}}_{t+1})$
where $\mathcal{T}_{s}$ and $\mathcal{O}^{{\textit{H}}}$ constitute the world model, while $\pi^{\textit{H}}$ and $\mathcal{T}_{z}$ are the human model. [^1] Since we have defined a general POMDP, it is convenient to assume that all parties’ predictive uncertainty about the world is part of their state distribution. This allows us to treat the human and the AI as sharing the same world model but possibly with different beliefs over the state. We can then express the expected utility of the human in the POMDP from any $({s},{z}^{\textit{H}})$.
$E\,{U}({s}_{k},{z}^{\textit{H}}_{k})\;=\;\mathbb{E}_{\tau_{k:k+N}\sim P(\cdot\mid s_{k},{z}^{\textit{H}}_{k})}\bigl{[}{U}(\tau_{k:k+N})\bigr{]}$ (2)
The key mathematical insight underpinning RLHS is that, for feedback purposes, the human’s internal state at the end of the interaction ${z}^{\textit{H}}_{k}$ should only affect the predicted rollout distribution through the human’s behavior ${\pi}^{\textit{H}}(\cdot|{z}^{\textit{H}}_{t})$; it should not contaminate the prediction of any given action’s outcome. By examining the below definition, we can see that this condition is readily violated by the value approximated by immediate feedback.
@@
The human’s foresight value at time $k\geq 0$ is the utility expected from $k$ to $k+N$ under the human’s subjective prediction of future outcomes:
$\overrightarrow{{V}}_{k+N}({z}^{\textit{H}}_{k})\;:=\;\mathbb{E}_{s_{k}\sim b^{\textit{H}}_{k}(\cdot|{z}^{\textit{H}}_{k})}\bigl{[}EU(s_{k},{z}^{\textit{H}}_{k})\bigr{]}$ (3)
where $b_{k}^{\textit{H}}(\cdot\mid{z}^{\textit{H}}_{k})$ is the human’s belief over the state, informed by their internal state after the interaction.
The above expectation is taken over the human’s belief. Whether we consider the belief of the user or a separate evaluator, it is a function of the internal state ${z}^{\textit{H}}_{k}$ resulting from the interaction, and therefore it can, in general, be influenced by the AI output.
Definition 4 (Expected Hindsight Value).
The AI’s expected hindsight value at time $k\geq 0$ is the expected reward-to-go from $k$ to $k+N$ under the AI’s world model.
$\overleftarrow{{V}}_{k+N}({z}^{\textit{H}}_{k})\;:=\;\mathbb{E}_{s_{k}\sim P^{W}(\cdot|z_{k}^{W})}\bigl{[}EU(s_{k},{z}^{\textit{H}}_{k})\bigr{]}$ (4)
@@
Unlike the foresight value, the expected hindsight value relies on the AI world model’s internal state, which can be kept independent from the AI interaction output (e.g. in the case of a LLM world model, by excluding the interaction from its prompt). Our proposed RLHS scheme performs stochastic gradient ascent on this expectation by repeatedly sampling world states ${s}_{k}$ and rolling out the interaction outcome. Fig. 2 illustrates the difference between foresight and hindsight feedback. In Appendix B, we show theoretically that providing human evaluators with hindsight during RLHF generally reduces misalignment and improves utility.
Figure 2: Illustration of the difference between RLHF and RLHS: In RLHF, the evaluator must predict the future outcome, whereas in RLHS, the outcome is sampled from the AI’s world model, independent of the AI interaction output.
@@ 3.2 Alignment with hindsight simulation
Hindsight Simulation. To translate our theoretical insights into practical implementation, we introduce the concept of hindsight simulation—the cornerstone of our RLHS framework. Hindsight simulation allows evaluators, whether human or AI, to make more informed decisions based on simulated outcomes. In practice, hindsight simulation can involve collecting feedback from human evaluators or employing another language model as a proxy to simulate the feedback process. After a real or simulated user takes an action based on AI suggestions (e.g., purchasing an item), the world model simulates the outcome (e.g., whether the purchased item meets the desired criteria). The evaluator then provides feedback informed by both the simulated outcome and their prior interaction with the AI model.
@@
Illustrative Example: Marketplace Chatbot. We demonstrate the practical impact of RLHS by fine-tuning various open-source LLMs on a marketplace consultancy task. The chatbot’s goal is to assist customers in making purchasing decisions by providing recommendations based on available product information. We assume that both customers and the chatbot have access to some public information, such as product lists and prices, but customers have their internal preferences, e.g., wanting a TV with 8K resolution, that are unknown to the chatbot. To our knowledge, existing RLHF schemes deployed for training marketplace chatbots (e.g., Amazon, 2024) use customer feedback based on the interaction (i.e., satisfaction on the spot) but not on its downstream outcome (i.e., whether the purchased item actually meets their preferences), which may cause emergent misalignment by incentivizing the chatbot to manipulate user predictions.
Hindsight simulation should mitigate this issue by decoupling feedback from outcome prediction. We simulate customers who interact with the chatbot, make a purchase decision, observe the outcome, and give feedback based on overall satisfaction—and compare with immediate feedback.

### v3

@@ Abstract
While Reinforcement Learning from Human Feedback (RLHF) has shown promise in aligning generative AI, we present empirical evidence that it can also cause severe, systematic misalignment. We hypothesize that this stems from evaluator feedback depending on downstream outcome predictions (foresight) that can be influenced by the AI’s output, inducing Goodhart’s law dynamics. We present a theoretical analysis showing that conditioning evaluator feedback on downstream observations (hindsight) inhibits this effect by decoupling the alignment signal from potentially compromised predictions—crucially, the result holds even if the observed outcomes are sampled from the AI’s own world model. Building on this insight, we introduce Reinforcement Learning from Hindsight Simulation (RLHS), which presents plausible simulated outcomes to evaluators before eliciting feedback. We validate RLHS across three consultancy settings—marketplace interactions, restaurant recommendations, and online course advising—using both online (PPO) and offline (DPO) fine-tuning methods, and show that it substantially improves alignment over RLHF in experiments and human evaluations. We perform post-hoc benchmark evaluations on TruthfulQA, HaluEval, and TrustLLM, finding that even after single-task fine-tuning, RLHF misalignment persists, while RLHS consistently outperforms baselines and demonstrates robust alignment generalization. The project webpage and code are available at https://rl-hindsight.github.io.
@@ 1 Introduction
Aligning artificial intelligence (AI) systems with human values and goals is crucial to ensuring their behavior is helpful, honest, and trustworthy. Eliciting human feedback is a widely-used alignment strategy (Leike et al., 2018), with successful applications to, e.g., training AI assistants (Glaese et al., 2022; Touvron et al., 2023; Anthropic, 2023; Achiam et al., 2023). In particular, Reinforcement Learning from Human Feedback (RLHF) (Christiano et al., 2017; Ziegler et al., 2019; Ouyang et al., 2022; Stiennon et al., 2020) leverages human feedback to fine-tune and align foundation models (FMs). While RLHF has shown promise in aligning models with human preferences, it relies predominantly on immediate assessments of isolated interactions, which may not accurately account for their downstream outcomes. Inaccurate user or evaluator feedback can misguide the model’s behavior and undermine the alignment process  (Casper et al., 2023; Pandey et al., 2022; Chmielewski & Kucker, 2020). On the other hand, recent theoretical work suggested that feedback from users who cannot fully observe an AI assistant’s actions could lead RLHF to learn deceptive behaviors (Lang et al., 2024).
@@
Figure 1: RLHF can incentivize AI systems to provide deceptive or inaccurate information by prioritizing immediate positive feedback over long-term outcomes. For example, a customer may initially prefer optimistic shopping advice but ultimately regret an ill-informed purchase. The proposed RLHS method removes this incentive by simulating downstream outcomes before eliciting feedback. The resulting fine-tuned AI models show superior alignment with users’ underlying utility.
@@
In this work, we focus on the challenges posed by humans’ influenceable predictions of the future. In many settings, the utility provided by an AI system to a human user (and similarly its “helpfulness” and “harmlessness”, which RLHF evaluators are typically asked to assess) is not an intrinsic property of the outputs that it generates but rather a function of their real-world consequences, brought about by the user’s real-world decisions upon consuming said outputs. Our central insight is that rewarding an AI system to improve users’ or evaluators’ in-the-moment assessments of interactions creates a pernicious Goodhart’s law dynamic: it incentivizes the AI system to favor outputs that induce unrealistically optimistic expectations in users. While at best these may be innocuous, at worst they can lead users to make poor choices resulting in degraded or even unsafe outcomes.
@@
We provide substantial empirical evidence that indeed this phenomenon can arise even in simple settings: we find that immediate human feedback elicited at the end of the human–AI interaction frequently misrepresents true utility in consultancy-type interactions, and, when used as a proxy for it in RLHF fine-tuning, it systematically drives misalignment with human goals (Fig. 1, top). Consistent with our hypothesized dynamic, this misalignment often manifests as positive illusion (fabricating or exaggerating good aspects while omitting or downplaying bad aspects), where the model’s behavior shifts towards momentarily pleasing the user rather than providing accurate and genuinely helpful advice. This systematically leads users to make ill-informed decisions whose poor downstream outcomes contrast starkly with their high satisfaction rating at the end of the interaction.
@@
To address these challenges, we propose a simple but effective misalignment mitigation mechanism: letting evaluators experience the downstream outcomes of each interaction before gathering their feedback. To circumvent the material and ethical difficulties in exposing real people to real consequences, we introduce a novel alignment fine-tuning methodology called Reinforcement Learning from Hindsight Simulation (RLHS), which uses a pretrained world model to simulate likely human decisions and their downstream outcomes after each generated output and presents these to evaluators for feedback. Our key finding is that granting evaluators the benefit of hindsight—and relieving them of the burden of foresight—significantly reduces model misalignment after fine-tuning: even if the AI’s own world model contains inaccuracies, these are independent of the outputs presented to the user, and therefore the AI has no incentive to distort them.
We evaluate hindsight simulation with both offline and online preference optimization approaches, including direct preference optimization (DPO) (Rafailov et al., 2024) and proximal policy optimization (PPO) (Schulman et al., 2017) and find that it greatly improves alignment in both paradigms.
Figure 2: Comparison between RLHF and RLHS: In RLHF, the evaluator predicts the future outcome, while RLHS samples it from the AI’s world model, independent of the AI interaction output.
We further validate these results through an online user study, where RLHS consistently improved objective utility and subjective satisfaction of our participants. Our comparative findings demonstrate that RLHS outperforms non-hindsight methods—specifically Reinforcement Learning from AI Feedback (RLAIF), which similarly uses AI generation as a proxy for human feedback, and has been shown to produce results similar to RLHF (Bai et al., 2022b; Lee et al., 2023). Finally, we evaluate our fine-tuned models on three benchmarks: TruthfulQA (Lin et al., 2021), HaluEval (Li et al., 2023a), and TrustLLM (Sun et al., 2024), covering hallucination, sycophancy, and privacy. Results show that RLHS consistently outperforms baselines and demonstrates strong out-of-domain generalization.
@@ 2 Algorithm: RL from Hindsight Simulation
We begin by examining how RLHF can become misaligned due to the limitations of human foresight. We then discuss the advantages of incorporating hindsight and explain why and how it mitigates misalignment. Finally, we introduce RLHS—a principled alignment method that employs simulated hindsight to decouple feedback from potentially AI-compromised predictions. We provide relevant background and formal definitions in Appendix C.
@@ 2.1 Hindsight Mitigates Misalignment
We consider the setting in which the human user interacts with an AI up to time $k$ (interaction phase), and then proceeds to take actions (acting phase). We are interested in the user’s acting phase utility. Let $N$ be a sufficiently long horizon for the acting phase outcomes to be experienced by the user.
Definition 1 (Human utility).
Let $\tau_{k:k+N}=({s}_{t},{a}^{\textit{H}}_{t})_{t=k}^{k+N}$ denote the trajectory of world states and human actions from time $k\geq 0$ to $k+N$. The user’s utility is the discounted sum of rewards, parametrized by human preferences ${\theta}^{\textit{H}}$ (unknown to the AI a priori):
@@
Definition 2 (Human trajectory distribution).
Let $P(\tau_{k:k+N}\mid s_{k},z^{H}_{k})$ be the probability distribution over the trajectory $\tau_{k:k+N}=({s}_{t},{a}^{\textit{H}}_{t})_{t=k}^{k+N}$ induced from initial conditions $({s}_{k},{z}^{\textit{H}}_{k})$ by the human user’s observations ${o}^{\textit{H}}_{t}$, internal states ${z}^{\textit{H}}_{t}$, and actions ${a}_{t}^{\textit{H}}$:
@@
${o}^{\textit{H}}_{t}$ $\sim\mathcal{O}^{\textit{H}}(\cdot\mid{s}_{t}),$ $\quad z_{t+1}^{H}$ $\sim\mathcal{T}_{z}(\cdot\mid{z}^{\textit{H}}_{t},{a}_{t}^{\textit{H}},{o}^{\textit{H}}_{t+1})$
where $\mathcal{T}_{s}$ and $\mathcal{O}^{{\textit{H}}}$ constitute the world model, while $\pi^{\textit{H}}$ and $\mathcal{T}_{z}$ comprise the human behavior model. [^1] Since we have defined a general partially observable Markov decision process (POMDP), it is convenient to assume that all parties’ predictive uncertainty about the world is part of their state distribution. This allows us to treat the human and the AI as sharing the same world model but possibly with different beliefs over the state. We can then express the expected utility of the human from any $({s}_{k},{z}^{\textit{H}}_{k})$.
$E\,{U}({s}_{k},{z}^{\textit{H}}_{k};{\theta}^{\textit{H}})\;=\;\mathbb{E}_{\tau_{k:k+N}\sim P(\cdot\mid s_{k},{z}^{\textit{H}}_{k})}\bigl{[}{U}(\tau_{k:k+N};{\theta}^{\textit{H}})\bigr{]}$ (2)
The key mathematical insight underpinning RLHS is that, for feedback purposes, the only effect of the human’s post-interaction internal state ${z}^{\textit{H}}_{k}$ on the predicted trajectory distribution should be through the human’s subsequent behavior ${\pi}^{\textit{H}}(\cdot|{z}^{\textit{H}}_{t})$; that is, ${z}^{\textit{H}}_{k}$ should not contaminate the prediction of any given action’s outcome. By examining the below definition, we can see that this condition is readily violated by when the user’s value after an interaction is approximated by immediate feedback.
@@
The human’s foresight value, for any underlying preferences ${\theta}^{\textit{H}}$, is the utility expected under the human’s subjective prediction of future outcomes:
$\overrightarrow{{V}}_{k+N}({z}^{\textit{H}}_{k};{\theta}^{\textit{H}})\;:=\;\mathbb{E}_{s_{k}\sim b^{\textit{H}}_{k}(\cdot|{z}^{\textit{H}}_{k})}\bigl{[}EU(s_{k},{z}^{\textit{H}}_{k};{\theta}^{\textit{H}})\bigr{]}$ (3)
where $b_{k}^{\textit{H}}(\cdot\mid{z}^{\textit{H}}_{k})$ is the human’s belief over the state, informed by their internal state after interaction.
The above expectation is taken over the human’s belief. Whether we consider the belief of the user or a separate evaluator, it is a function of the internal state ${z}^{\textit{H}}_{k}$ resulting from the interaction, and therefore it can, in general, be influenced by the AI output—opening a path to reward hacking.
Definition 4 (AI-expected hindsight value).
The AI-expected hindsight value at time $k\geq 0$, for any underlying user preferences ${\theta}^{\textit{H}}$, is the expectation of the user’s utility under the AI’s world model.
$\overleftarrow{{V}}_{k+N}({z}^{\textit{H}}_{k};{\theta}^{\textit{H}})\;:=\;\mathbb{E}_{s_{k}\sim P^{W}(\cdot|z_{k}^{W})}\bigl{[}EU(s_{k},{z}^{\textit{H}}_{k};{\theta}^{\textit{H}})\bigr{]}$ (4)
@@
Unlike the foresight value, the expected hindsight value relies on the AI world model’s internal state, which can be kept independent from the AI interaction output (e.g. in the case of a LLM world model, by excluding the interaction from its prompt). Our proposed RLHS scheme performs stochastic gradient ascent on this expectation by repeatedly sampling world states ${s}_{k}$ and rolling out the interaction outcome. Fig. 2 illustrates the difference between foresight and hindsight feedback. In Appendix D, we show theoretically that providing human evaluators with hindsight during RLHF generally reduces misalignment and improves utility.
@@ 2.2 Alignment with hindsight simulation
Hindsight Simulation. To translate our theoretical insights into practical implementation, we introduce the concept of hindsight simulation—the cornerstone of our RLHS framework. Hindsight simulation allows evaluators, whether human or AI, to make more informed decisions based on simulated outcomes. In practice, hindsight simulation can involve feedback from human evaluators or using another language model as a proxy. After a real or simulated user takes an action based on AI suggestions (e.g., purchasing an item), the world model simulates the outcome (e.g., whether the purchased item meets the desired criteria). The evaluator then provides feedback informed by both the simulated outcome and their prior interaction with the AI model.
@@
Illustrative Examples: Consultancy Chatbot. We demonstrate the practical impact of RLHS by fine-tuning various open-source LLMs on consultancy chatbot tasks. The chatbot’s goal is to assist users in making decisions by providing recommendations based on available information. We assume that both users and the chatbot have access to some public information, but users have internal preferences unknown to the chatbot. To our knowledge, existing RLHF schemes deployed for training consultancy chatbots (e.g., Amazon, 2024) use user feedback based on the interaction (i.e., satisfaction on the spot) but not on its downstream outcome (i.e., whether the decision actually meets their preferences), which may cause emergent misalignment by incentivizing the chatbot to manipulate user predictions. Hindsight simulation should mitigate this issue by decoupling feedback from outcome prediction. Specifically, we simulate scenarios where users interact with chatbots, make decisions, observe outcomes, and subsequently provide feedback based on overall satisfaction. We compare this approach against immediate feedback mechanisms.